OpenStudy (dakid88):

Arnob, Bella, Colin, Dante, and Erin are going to a baseball game. They have a total budget of $100.00 to spend. Each game ticket costs $17.50 and each drink costs $2.00. The inequality below relates x, the number of drinks the 5 friends could buy in all, with their ticket costs and budget. Which best describes the restrictions on the number of drinks they can buy? They can buy only from 0 to 3 drinks, but no more. They can buy only from 0 to 6 drinks, but no more. They can buy only from 0 to 7 drinks, but no more. They can buy only from 0 to 12 drinks, but no more.

OpenStudy (dakid88):

the inequality is this\[5(17.5)+2x \le100\]

OpenStudy (babylove18):

17.50 * 5 =87.50 100-87.50=12.50 12.50/2=6.25 so b

OpenStudy (dakid88):

thxs